The mole fraction of the solute in one molar aqueous solution is ________.

Grade:11

1 Answers

Umakant biswal
7 years ago
1  molal aqueous solution means 1 mole of solute dissolved in 1 kg of water . 
no of moles of solute = 1 
no of  moles of water = 1000/18 = 55.556 
mole fraction of solute = no of moles of solute / no of moles of solute + solvent 
mole fraction 
= 1/ 1+ 55.556 
= 0.0177 
HOPE IT CLEARS 
ALL THE BEST ..